OpenStudy (anonymous):

Find the quadratic function y=ax^2+bx+c whose graph passes through the given points: (1,1),(3,-17),(-2,13)

OpenStudy (anonymous):

i am not sure but maybe you can solve for this system of equations 1=a+b+c -17=a3^2+b3+c 13=a*-2^2+b*-2+c

OpenStudy (anonymous):

i really do appreciate you hun :) i've been working on this problem for over 30 minutes before I decided to post it

OpenStudy (anonymous):

? how do you solve it from here...I am so lost

OpenStudy (anonymous):

well this one is messy. solve a in eq 1 plug it into eq 2 solve for b in eq 2 plug it into eq 3

OpenStudy (anonymous):

a=-1 b=-5 c = 7

OpenStudy (anonymous):

y=ax^2+bx+c but its supposed to look like this... y=-x^2-5x+7?

OpenStudy (anonymous):

y=ax^2+bx+c y=-1*x^2-5*x+7 so yes that is the answer :D

OpenStudy (anonymous):

:D you are a life saver hun!!

OpenStudy (anonymous):

:D not yet see if it is right Hope it is :)

OpenStudy (anonymous):

i really like you, you're so helpful! :) it was right, its an online class with immediate feedback haha

OpenStudy (anonymous):

well so long as you know how that got done :)

OpenStudy (anonymous):

um...what if the numbers change... (-1,-8),(1,-2),(-3,2)

OpenStudy (anonymous):

you lost me...sorry :/

OpenStudy (anonymous):

make a system of equations again -8=a*(-8)^2+b*(-1)+c, -2=a+b+c, 2=a*(-3)^2+b*(-3)+c

OpenStudy (anonymous):

they're fractions.... ? now i'm really confused

OpenStudy (anonymous):

it doesn't matter just plug them in :)

OpenStudy (anonymous):

ha, can i ask how old you are....you're helping me with college algebra and this class is constantly going over my head...

OpenStudy (anonymous):

y=(-8/59)*x^2+(-75/59)*x+(-35/59)

OpenStudy (anonymous):

I am too old :P lol

OpenStudy (anonymous):

26

OpenStudy (anonymous):

[-8=a*(-1)^2+b*(-1)+c, -2=a+b+c, 2=a*(-3)^2+b*(-3)+c]...one number off...It kept saying it was incorrect...i understand where you put the numbers and why!!...ok, maybe I'm too excited about that but *sigh*, you up for another one :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how many chances you get to get it right :P

OpenStudy (anonymous):

3...i had already used one and then i entered the fractions so i had to get it right the third time, you definitely got me closer then what i had in the first try :)

OpenStudy (anonymous):

was y=(-8/59)*x^2+(-75/59)*x+(-35/59) correct?

OpenStudy (anonymous):

sometimes computers are picky -(8/59)*x^2-(75/59)*x-35/59 that might work Unless i did something completly wrong

OpenStudy (anonymous):

[-8=a*(-1)^2+b*(-1)+c, -2=a+b+c, 2=a*(-3)^2+b*(-3)+c] this one was...you had put -8 here ^

OpenStudy (anonymous):

ha oops the first -1

OpenStudy (anonymous):

oh yes sorry i made mistake there

OpenStudy (anonymous):

its not a problem, that is the ONLY mistake you made...in my book, that's pretty dang good hun :)

OpenStudy (anonymous):

(-1,-8),(1,-2),(-3,2) -8=a*(-1)+b*(-1)+c, -2=a+b+c, 2=(-3)*a+(-3)*b+c

OpenStudy (anonymous):

still getting error grr lol

OpenStudy (anonymous):

did you put the [ ] thingys around it?

OpenStudy (anonymous):

yes when i paste to the website

OpenStudy (anonymous):

ooh i know :)

OpenStudy (anonymous):

forgot to square :P

OpenStudy (anonymous):

and it still gives you an error...

OpenStudy (anonymous):

ha glad you figured it out...we'd be here all night if you were counting on me finding the error...sometimes i hate being dyslexic

OpenStudy (anonymous):

(-1,-8),(1,-2),(-3,2) -8=a*(-1)^2+b*(-1)+c, -2=a+b+c, 2=(-3)^2*a+(-3)*b+c

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

oh men I need glasses not the smae but it sucks :P looking at a lot of numbers and messy equations :P

OpenStudy (anonymous):

now the solution looks cute :D by that i mean nice integers :P

OpenStudy (anonymous):

I have glasses too...i just never wear them...nicely done hun!

OpenStudy (anonymous):

and I mean same not smae :P

OpenStudy (anonymous):

its all good, in some ways being dyslexic helps me decode things too :) hahaha

OpenStudy (anonymous):

I am myopic

OpenStudy (anonymous):

that is awesome :) I suck at spelling. Maybe my fingers are dyslexic.

OpenStudy (anonymous):

I always switch letters. maybe it has to do more with me being mexican and that I try to spell things in spanish at times.

OpenStudy (anonymous):

haha, i never thought of that, i dont know what my vision is called, i can see about to the hood of my car...and the rest is really blurry...nearsidedness ha :)

OpenStudy (anonymous):

myopic is near sighted :) idk what far sighted is lol.

OpenStudy (anonymous):

oh...haha wow, i feel smarter already :P <--this sir..is a legit face

OpenStudy (anonymous):

This website will make you feel smarter :P

OpenStudy (anonymous):

ha fer sure... :)

OpenStudy (anonymous):

better than FB :P

OpenStudy (anonymous):

fer sure, you throw a ball straight up from a rooftop. the ball misses the rooftop on its way down and eventually strikes the ground. a mathematical model can be used to describe the relationship for the ball's height above the ground, y, after x, seconds.